# Optimizing EPR pulses for broadband excitation and refocusing

### v1

@@ Abstract
In this paper, we numerically optimize broadband pulse shapes that maximize Hahn echo amplitudes. Pulses are parameterized as neural networks (NN), nonlinear amplitude limited Fourier series (FS), and discrete time series (DT). These are compared to an optimized choice of the conventional hyperbolic secant (HS) pulse shape. A power constraint is included, as are realistic shape distortions due to power amplifier nonlinearity and the transfer function of the microwave resonator. We find that the NN, FS, and DT parameterizations perform equivalently, offer improvements over the best HS pulses, and contain a large number of equivalent optimal solutions, implying the flexibility to include further constraints or optimization goals in future designs.
@@ I Introduction
The use of shaped pulses in electron paramagnetic resonance (EPR) spectroscopy is a topic of recent interest Prisner et al. (2001); Doll et al. (2013); Spindler et al. (2017); Doll and Jeschke (2017); Tait and Stoll (2017); Prisner (2019). They address the basic challenge that the excitation bandwidth of monochromatic square pulses is much smaller than the spectral line width of samples. This situation can arise in nuclear magnetic resonance (NMR), but it is more broadly relevant in EPR. Typical EPR spectral widths are about 250 MHz for a nitroxide at Q-band ($\approx$ 1.2 T) or $\approx$ 2 GHz for a Cu(II) complex at X-band ($\approx$ 0.35 T). Broadband excitation using shaped pulses increases sensitivity and mitigates undesirable effects like orientation selection.
@@
However, HS pulses are still not truly optimal. In principle, pulses should be as short as possible in order to minimize relaxation effects, but in practice, limited available power limits the maximum achievable pulse amplitude which in turn limits the frequency sweep rate and therefore the pulse length. In response, optimal control methods such as composite pulses Counsell et al. (1985); Levitt (1986), adiabatic pulses Baum et al. (1985); Garwood and DelaBarre (2001), optimal control theory (OCT) pulses using different numerical algorithms Cano et al. (2002); Skinner et al. (2003); Kobzar et al. (2004); Foroozandeh et al. (2019); Haller et al. (2022), including gradient ascent pulse engineering (GRAPE) Khaneja et al. (2005); Nielsen et al. (2010); Tošner et al. (2009); Borneman and Cory (2012); Goodwin and Kuprov (2016); Spindler et al. (2012), were developed to accomplish broadband excitation and uniform inversion across a given bandwidth. In practice, optimal shaped pulses are distorted by nonlinearities in the power amplifier and by the resonator transfer function, moving them away from the extremum in the optimization landscape. One can try to compensate for these distortions post-optimization, but the necessary compensation may not be possible while respecting constraints such as limited power.
In this paper, we use an optimization method that includes a model of the full experimental transfer chain and resulting shape distortions. We investigate the use of variously parameterized broadband pulses for a Hahn echo sequence and the effect of varying the pulse length ratio on echo amplitude, refocusing time, and refocusing phase. Performances of individual pulses as well as of the entire pulse sequence are compared. Section II summarizes the transmitter model and the spin physics model, section III describes the pulse parameterizations used, section IV provides details about the optimization method, and section V discusses the results.
@@ II Model
These equations only model amplitude-to-amplitude modulation (AM/AM) effects of the amplifier and neglect possible amplitude-to-phase modulation (AM/PM) effects. AM/PM effects would alter the term $\phi(t)$ in Eq. (7) by mixing the I and Q signals.
@@
Note that there is a subtle difference between imposing separate constraints on $I$ and $Q$ as we do here versus simply imposing an overall amplitude constraint. Constraining $I$ and $Q$ independently restricts the accessible domain in the I/Q plane to a square rather than a circle. Maximum power is only attained at the four corners of the square when $|I|=|Q|=1$, i.e., phases of ±$\pm 45^{\circ},\pm 135^{\circ}$. Changing the phase of a maximum-power pulse with one of these phases will result in a power loss.
@@ III Parameterizations
We start by considering the original KBB broadband pulse sequence using generalized HS shapes for the $\pi/2$ and $\pi$ pulses. Then, while retaining the general bipartite structure of a $\pi/2$ pulse followed by a $\pi$ pulse, to introduce more shape flexibility, we consider three models with significantly more parameters: a nonlinear amplitude-limited Fourier series (FS), a discrete-time series (DT), and a neural network (NN). We do not constrain the pulse flip angles to be $\pi/2$ and $\pi$, but we will still refer to them by those labels in continuity with the KBB design and in anticipation that the optimization process will indeed drive them to be such.
@@ IV Optimization
For the FS parameterization, we use $24$ terms per pulse shape, totaling $96$ free parameters for the sequence. For the DT parameterization, we use a time step of $\Delta t=0.625$ ns, resulting in $384$ parameters for $T_{\pi/2}+T_{\pi}=$ 120 ns. The NN parameterization uses $644$ free parameters, as described in Sec. III.4. We used the Julia package DiffEqFlux.jl Rackauckas et al. (2021) to form the NN parameterization and the BFGS optimizer from the Zygote.jl and Optim.jl packages for optimizing the various pulse shape parameterizations Innes (2019). The limiting factor in computational cost is the numerical solution of Schrödinger’s equation for each of the $250$ Larmor frequencies for both pulses at each optimization step. We solved these in parallel on cluster computing resources with the Bogacki–Shampine method (BS5) as implemented in the DifferentialEquations.jl package. All optimizations were terminated upon the condition that the difference between the current objective function value and the objective function value $20$ steps previous was less than $10^{-4}$. All three large parameterizations took around $500$-$1000$ optimization steps to converge, whereas the 5-parameter hyperbolic secant model only took tens of steps.
Figure 4: The optimized pulse shapes as seen by the spins after passing through the chain of transfer functions shown in Fig. 1, along with the formed echoes. The optimized HS $\pi/2$ and $\pi$ pulse lengths are $80$ and $40$ ns long respectively while the FS, DT, and NN have $\pi/2$ and $\pi$ pulse lengths of $60$ ns each. The phase, $\phi$, of each echo is also reported.
Figure 5: Results of 20 optimizations each for the three large parameterizations of Fourier series (FS), discrete time (DT), and neural network (NN) with $80/40$ ns, $/60$ 60 ns, and $40/80$ ns pulse lengths from random starting points. The top panel shows the echo amplitudes for the different pulse types, where the dashed lines are the echo amplitudes of the best performing hyperbolic secant (HS) pulses for each respective pulse length ratio. The error bars and dotted lines show the effect of altering the amplifier compression function as described in the text. The bottom three panels show the relative echo amplitude reduction for the case where $B_{1}$ field is reduced to $80\%$ of its original value, the time between the end of the $\pi$ pulse and the echo maximum, and the phase with which the spins are refocusing, all versus the echo amplitudes.
@@ V Results
We performed the optimizations using the parameter values shown in Table 1, representative of a Q-band EPR spectrometer with a nitroxide sample. The carrier frequency of $33.65$ GHz corresponds to a static magnetic field strength of about $1.2$ T. The value of $\varGamma$ corresponds to a $3$-dB bandwidth of $450$ MHz for the AWG. The loaded Q-value corresponds to a resonator $3$-dB bandwidth of $168$ MHz. The maximum power limit corresponds to an oscillatory magnetic field strength of about $3$ mT. The choice of a 120 ns total pulse time was made in order to examine a case where the power constraint starts to deteriorate the performance of the hyperbolic secant pulse. The delay time, $\tau_{1}$, was chosen to be $100$ ns. While this value may affect the path the optimizer takes through the optimization landscape, pulse shapes optimized with one value of $\tau_{1}$ produce the same echo amplitude for a different value of $\tau_{1}$. The nonlinear phase dispersion of the $\pi/2$ pulse will still be cancelled by that of the $\pi$ pulse, and changing the delay time between them only changes the linear part of the phase dispersion, thus changing the echo time and phase but not the echo amplitude.
@@
The best HS result overall was for a 2:1 pulse length ratio, with $A_{\pi/2}/2\pi=23.42$ MHz and $A_{\pi}/2\pi=61.66$ MHz for the amplitudes of the $\pi/2$ and $\pi$ pulses, $\beta=7.09$, $n=1.57$, and $\Delta\omega_{\mathrm{BW}}/2\pi=138.51$ MHz, which produces an echo amplitude of $0.9070$ (see Fig. 4 top). To quantify the effect of artificially allowing the HS to exceed the power limit as discussed in Sec. IV, we also optimized strictly within the inscribing circle of the allowed domain in the I/Q plane, $I(t)^{2}+Q(t)^{2}\leq 1$. This more restricted optimization resulted in an echo amplitude of $0.8621$ for our example parameters. Of course, the total extra power gained from accessing the corners of the square domain diminishes as the value $V_{\mathrm{sat}}$ decreases, since the amplifier saturates more readily. Having quantified this effect, below we set it aside and focus on the performance of other parameterizations compared to the HS with equivalent maximum power whose echo amplitude is $0.9070$.
For those other pulse parameterizations, all three pulse length ratios do allow for both full rotation across the bandwidth and refocusing, with a 1:1 pulse length ratio performing the best. The maximized echo amplitudes are, in increasing order, $0.9852$ for the FS, $0.9928$ for the DT, and $0.9951$ for the NN. These pulses plotted in Fig. 4 are the drive functions seen by the spins, i.e. after passing through the low-pass filter, the amplifier, and the resonator transfer function (thus the pulses are several ns longer than the nominal $60$ ns due to the finite ring-down time). The pulses show irregular shapes; there are no apparent interpretable features. Clearly, the FS, DT, and NN pulses offer performance gains compared to the HS for this scenario of short pulse time and limited power. Among these three parameterizations, though, there is not one that stands out as particularly advantageous. All are equivalently efficient computationally, taking around the same number of optimization steps.
@@
The error bars in Fig. 5A show the effect on the echo amplitude of reducing the amplifier compression $V_{\mathrm{sat}}$ by 60%, i.e. increasing amplifier compression without changing maximum output power. The FS, DT, and NN pulses are all relatively robust against this. This is because many of the numerically shaped pulses use as much power as possible, so only the maximum power of the amplifier matters rather than the shape of the compression function at intermediate power. This is not entirely clear from the plots of the output pulse shapes in Fig. 4, but the input pulse shapes of $I$ and $Q$ are typically toggling between their maximum values of $-1$ and $1$, and the values plotted are only diminished from maximum amplitude due to the effects of the low-pass filter and resonator. Thus, changing the general shape of the amplifying function does not affect the pulse shapes much. The HS pulse shapes, on the other hand, do use the full range of the amplifier, and the dotted lines show how their performance is substantially diminished under the same $V_{\mathrm{sat}}$ reduction.
In the $B_{1}$ field attenuation plot in Fig. 5B, we plot the fractional reduction of the echo amplitude when the $B_{1}$ field is reduced to $80\%$. This probes how robust the echo amplitude is to $B_{1}$ field inhomogeneity. The best HS pulse is one of the more robust pulses against $B_{1}$ field inhomogeneity, since it is constructed as an adiabatic frequency sweep. Other pulse types include equivalently robust pulses, but generally there is no correlation between echo amplitude and robustness to inhomogeneity because the pulses were not optimized for robustness.
@@
This ensemble of optimizations from different random seeds shows that there is no unique pulse shape that is optimal. Of course, there is always a trivial degeneracy of rotating both $I$ and $Q$ by ±$\pm 90^{\circ},\pm 180^{\circ}$ resulting only in a change of echo phase by the same angle (the degeneracy is four-fold rather than continuous due to the power constraint being a square rather than a circle in the I/Q plane), retaining the same refocusing time and echo amplitude. But beyond that, Fig. 5 shows that there are many pulse shapes with equivalent echo amplitudes that are not simply related by a phase transformation, as is clear from the different echo refocusing times. The fact that the landscape harbors many comparable minima suggests that there remains a significant amount of flexibility in the parameterized pulses that could be used to satisfy additional constraints, such as favoring a particular refocusing time or phase (as in Kallies and Glaser (2018)), robustness to $B_{1}$ inhomogeneity, echo phase independence from $\tau_{1}$, or some other desirable property.
@@
Figure 6 strongly suggests that the improvement of the optimal FS, DT, and NN pulses compared to the HS pulse comes from i) improved performance near the band edges and ii) improved intra-band $\pi$ rotation performance in the presence of the power constraint.
@@
In other words, since all the spin packet phases are flipped by the $\pi$ rotation, echo formation requires $\tau_{\mathrm{eff},\pi/2}(\omega)-\tau_{\mathrm{eff},\pi}(\omega)$ to be a constant Kallies and Glaser (2018). Figure 7 shows that this is indeed the case for the FS, DT, and NN pulses, which have nonlinear phase dispersions for the two pulses but cooperate such that their nonlinear parts mutually cancel. Note that the difference $\tau_{\mathrm{eff},\pi/2}(\omega)-\tau_{\mathrm{eff},\pi}(\omega)=t_{\mathrm{refocus}}-\tau_{1}$. For example, in this case, the FS pulse has $\tau_{\mathrm{eff},\pi/2}(\omega)-\tau_{\mathrm{eff},\pi}(\omega)=$-10 ns and so refocuses at $90$ ns while the NN pulse has $\tau_{\mathrm{eff},\pi/2}(\omega)-\tau_{\mathrm{eff},\pi}(\omega)=39$ ns and refocuses at $139$ ns.

### v2

@@ Abstract
In this paper, we numerically optimize broadband pulse shapes that maximize Hahn echo amplitudes. Pulses are parameterized as neural networks (NN), nonlinear amplitude limited Fourier series (FS), and discrete time series (DT). These are compared to an optimized choice of the conventional hyperbolic secant (HS) pulse shape. A power constraint is included, as are realistic shape distortions due to power amplifier nonlinearity and the transfer function of the microwave resonator. We find that the NN, FS, and DT parameterizations perform equivalently, offer improvements over the best HS pulses, and contain a large number of equivalent optimal maxima, implying the flexibility to include further constraints or optimization goals in future designs.
@@ I Introduction
The use of shaped pulses in electron paramagnetic resonance (EPR) spectroscopy is a topic of recent interest Prisner et al. (2001); Doll et al. (2013); Spindler et al. (2017); Doll and Jeschke (2017); Tait and Stoll (2017); Prisner (2019). They address the basic challenge that the excitation bandwidth of monochromatic square pulses is much smaller than the spectral line width of samples. This situation can arise in nuclear magnetic resonance (NMR), but it is more broadly relevant in EPR. Typical EPR spectral widths are about 250 MHz for a nitroxide at Q-band ($\approx$ 1.2 T) or $\approx$ 2 GHz for a Cu(II) complex at X-band ($\approx$ 0.35 T). Using shaped pulses can increase sensitivity and excitation bandwidth.
@@
However, HS pulses are not optimal. In practice, limited available power limits the maximum achievable pulse amplitude which in turn limits the frequency sweep rate and therefore puts a lower bound on the pulse time. However, relaxation can put an upper limit on pulse duration. Also, in some sequences, pulse lengths must be shorter than the evolution periods of the interactions of interest, for instance dipolar couplings in dipolar EPR. Endeward et al. (2023). Thus, along with the constraint of limited power, constraining the pulse time can cause the performance of pulses to suffer.
In response, optimal control methods such as composite pulses Counsell et al. (1985); Levitt (1986), adiabatic pulses Baum et al. (1985); Garwood and DelaBarre (2001), optimal control theory (OCT) pulses using different numerical algorithms Cano et al. (2002); Skinner et al. (2003); Kobzar et al. (2004); Foroozandeh et al. (2019); Haller et al. (2022), including gradient ascent pulse engineering (GRAPE) Khaneja et al. (2005); Nielsen et al. (2010); Tošner et al. (2009); Borneman and Cory (2012); Goodwin and Kuprov (2016); Spindler et al. (2012), were developed to accomplish broadband excitation and uniform inversion across a given bandwidth. In practice, optimal shaped pulses are distorted by nonlinearities in the power amplifier and by the resonator transfer function, moving them away from the extremum in the optimization landscape. One can try to compensate for these distortions post-optimization, but the necessary compensation may not be possible while respecting constraints such as limited power at fixed pulse time.
In this paper, we use an optimization method that includes a model of the full experimental transfer chain and resulting shape distortions while limiting both the available power along with the length of pulses. We investigate the use of variously parameterized broadband pulses for a Hahn echo sequence and the effect of varying the pulse length ratio on echo amplitude, refocusing time, and refocusing phase. Due to the freedom of the large parameter space and low number of constraints, we find the individual pulses ($\pi/2$ and $\pi$) act cooperatively as done previously with NMR COOP pulses Braun and Glaser (2010); Wei et al. (2014); Asami et al. (2018); Kallies and Glaser (2018). This shows that cooperatively performing pulses are still optimal under transmission distortions and in the presence of a power constraint while in limited pulse length time. Performances of individual pulses as well as of the entire pulse sequence are compared. Section II summarizes the transmitter model and the spin physics model, section III describes the pulse parameterizations used, section IV provides details about the optimization method, and section V discusses the results.
@@ II Model
These equations only model amplitude-to-amplitude modulation (AM/AM) effects of the amplifier and neglect possible amplitude-to-phase modulation (AM/PM) effects. AM/PM effects would alter the term $\varphi(t)$ in Eq. (7) by mixing the I and Q signals.
@@
Note that there is a subtle difference between imposing separate constraints on $I$ and $Q$ as we do here versus simply imposing an overall amplitude constraint. Constraining $I$ and $Q$ independently restricts the accessible domain in the I/Q plane to a square rather than a circle. Maximum power is only attained at the four corners of the square when $|I|=|Q|=1$, i.e., phases of ±$\pm 45^{\circ},\pm 135^{\circ}$. Changing the phase of a maximum-power pulse will generally result in a power loss. However, phase cycling can still be carried out in ±$\pm 90^{\circ}$ increments since this domain is symmetric under ±$\pm 90^{\circ}$ rotations.
@@ III Parameterizations
We start by considering the original KBB broadband pulse sequence using generalized HS shapes for the $\pi/2$ and $\pi$ pulses. KBB is not the only broadband refocusing sequence we could use. For instance, the CHORUS sequence Power et al. (2016a, b) uses linear swept pulses with effectively rectangular amplitude profiles that have more integrated power than hyperbolic secant pulses for a fixed $B_{1}$ amplitude and sequence time, and features improved robustness to $B_{1}$ field inhomogeneity Foroozandeh et al. (2019); Verstraete et al. (2021). However, we consider a strongly power-constrained regime with limited amplitude and pulse lengths where $B_{1}$ field inhomogeneity is a secondary concern. In this regime, HS pulses are preferable because they require less power off resonance, where it is costly to compensate for the profile of the resonator transfer function. Therefore, in this work we will use the KBB sequence with HS pulses as a baseline against which to compare other shaped pulses.
Then, while retaining the general bipartite structure of a $\pi/2$ pulse followed by a $\pi$ pulse, to introduce more shape flexibility, we consider three models with significantly more parameters: a nonlinear amplitude-limited Fourier series (FS), a discrete-time series (DT), and a neural network (NN). We do not constrain the pulse flip angles to be $\pi/2$ and $\pi$, but we will still refer to them by those labels in continuity with the KBB design and in anticipation that the optimization process will indeed drive them to be such.
@@ IV Optimization
For the FS parameterization, we use $24$ terms per pulse shape, totaling $96$ free parameters for the sequence. For the DT parameterization, we use a time step of $\Delta t=0.625$ ns, resulting in $384$ parameters for $T_{\pi/2}+T_{\pi}=$ 120 ns. The NN parameterization uses $644$ free parameters, as described in Sec. III.4. We used the Julia package DiffEqFlux.jl Rackauckas et al. (2021) to form the NN parameterization and the BFGS optimizer from the Zygote.jl and Optim.jl packages for optimizing the various pulse shape parameterizations Innes (2019). The limiting factor in computational cost is the numerical solution of Schrödinger’s equation for each of the $250$ Larmor frequencies for both pulses at each optimization step. We solved these in parallel on cluster computing resources with the Bogacki–Shampine method (BS5) as implemented in the DifferentialEquations.jl package. All optimizations were terminated upon the condition that the difference between the current objective function value and the objective function value $20$ steps previous was less than $10^{-4}$. All three large parameterizations took around $500$-$1000$ optimization steps to converge, whereas the 5-parameter hyperbolic secant model only took tens of steps. On average, these optimizations took around $2$-$3$ hours to complete running with $36$ cores in parallel. However, this substantial optimization time is not a problem, as it is a one-time computational cost for a given spectrometer setup. As long as the distortion chain has been properly characterized, the computed pulses should work without (or with minimal) spectrometer-based feedback optimization Goodwin et al. (2018). The robustness of these pulses to mischaracterization of the distortion chain elements and $B_{1}$ field inhomogeneity is discussed in Section V.
Figure 4: The optimized pulse shapes as seen by the spins after passing through the chain of transfer functions shown in Fig. 1, along with the formed echoes, the real part of the magnetization after rephasing the echo such that the maximum is completely real. The optimized HS $\pi/2$ and $\pi$ pulse lengths are $80$ and $40$ ns long respectively while the FS, DT, and NN have $\pi/2$ and $\pi$ pulse lengths of $60$ ns each. The phase, $\phi$, of each echo is also reported.
Figure 5: Results of 20 optimizations each for the three large parameterizations of Fourier series (FS), discrete time (DT), and neural network (NN) with $80/40$ ns, $/60$ 60 ns, and $40/80$ ns pulse lengths from random starting points. The top panel shows the echo amplitudes for the different pulse types, where the dashed lines are the echo amplitudes of the best performing hyperbolic secant (HS) pulses for each respective pulse length ratio. The $3$ circled points represent the $3$ best performing pulse shapes for each parameterization that are plotted in Fig. 4. The error bars and dotted lines show the effect of altering the amplifier compression function as described in the text. The bottom three panels show the relative echo amplitude reduction for the case where $B_{1}$ field is reduced to $80\%$ of its original value, the time between the end of the $\pi$ pulse and the echo maximum, and the phase with which the spins are refocusing, all versus the echo amplitudes.
@@ V Results
We performed the optimizations using the parameter values shown in Table 1, representative of a Q-band EPR spectrometer with a nitroxide sample. The carrier frequency of $33.65$ GHz corresponds to a static magnetic field strength of about $1.2$ T. The value of $\varGamma$ corresponds to a $3$-dB bandwidth of $450$ MHz for the AWG. The loaded Q-value corresponds to a resonator $3$-dB bandwidth of $168$ MHz. The maximum power limit of $\omega_{1,\mathrm{max}}=84$ MHz corresponds to an oscillatory magnetic field strength of about $3$ mT. The choice of a 120 ns total pulse time was made in order to examine a case where the power constraint starts to deteriorate the performance of the hyperbolic secant pulse. The delay time, $\tau_{1}$, was chosen to be $100$ ns. While this value may affect the path the optimizer takes through the optimization landscape, pulse shapes optimized with one value of $\tau_{1}$ produce the same echo amplitude for a different value of $\tau_{1}$. The nonlinear phase dispersion of the $\pi/2$ pulse will still be cancelled by that of the $\pi$ pulse, and changing the delay time between them only changes the linear part of the phase dispersion, thus changing the echo time and phase but not the echo amplitude.
The optimization and the resulting analysis does not consider any spins to be coupled. Chirped pulses in particular have been shown to create unwanted artifacts compared to rectangular pulses in situations with coupled electron and nuclear spins Doll and Jeschke (2016); Pribitzer et al. (2016). A separate analysis can be performed to determine the effect that these optimized pulse shapes have on multi-dimensional spectra.
@@
The best HS result overall was for a 2:1 pulse length ratio, with $A_{\pi/2}/2\pi=23.42$ MHz and $A_{\pi}/2\pi=61.66$ MHz for the amplitudes of the $\pi/2$ and $\pi$ pulses, $\beta=7.09$, $n=1.57$, and $\Delta\omega_{\mathrm{BW}}/2\pi=138.51$ MHz, which produces an echo amplitude of $0.9070$ (see Fig. 4 top). To quantify the effect of artificially allowing the HS to exceed the power limit as discussed in Sec. IV, we also optimized strictly within the inscribing circle of the allowed domain in the I/Q plane, $I(t)^{2}+Q(t)^{2}\leq 1$. This more restricted optimization resulted in an echo amplitude of $0.8621$ for our example parameters. Of course, the total extra power gained from accessing the corners of the square domain diminishes as the value $V_{\mathrm{sat}}$ decreases, since the amplifier saturates more readily. Having quantified this effect, below we set it aside and focus on the performance of other parameterizations compared to the HS with equivalent maximum power whose echo amplitude is $0.9070$. As an alternative point of comparison, the time length required for an HS pulse sequence to perform as well as the other optimized pulse shapes is $180$ ns, $50\%$ longer than the other pulses.
For those other pulse parameterizations, all three pulse length ratios do allow for both full rotation across the bandwidth and refocusing, with a 1:1 pulse length ratio performing the best. The maximized echo amplitudes are, in increasing order, $0.9852$ for the FS, $0.9928$ for the DT, and $0.9951$ for the NN. These pulses plotted in Fig. 4 are the drive functions seen by the spins, i.e. after passing through the low-pass filter, the amplifier, and the resonator transfer function (thus the pulses are several ns longer than the nominal $60$ ns due to the finite ring-down time). The pulses show irregular shapes; there are no apparent interpretable features. Clearly, the FS, DT, and NN pulses offer performance gains compared to the HS for this scenario of short pulse time and limited power. Among these three parameterizations, though, there is not one that stands out as particularly advantageous. As mentioned in Sec. III, the reasoning behind using the NN was to efficiently represent possible correlations between the two pulses. However, both the FS and DT show equivalent cooperativity in compensating for phase accumulation between the $\pi/2$ and $\pi$ shapes, and the NN parameterization offered no extra advantage in that regard. The pulse shapes before passing through the distortion chain are plotted in Fig. S1 of the Supplementary Material sup .
All three parameterizations are equivalently efficient computationally, taking around the same number of optimization steps. The reason for the wide range of performance with many equivalent maxima is that the landscape contains many local maxima. Any large parameterization flexible enough to access a large area of that landscape in an unbiased way will lead to many maxima of varying quality. The outcome of a particular optimization depends on the initial point, so it is clearly useful to try many random initial seeds as we have done.
@@
The error bars in Fig. 5A show the effect on the echo amplitude of reducing the amplifier compression $V_{\mathrm{sat}}$ by 60%, i.e. increasing amplifier compression without changing maximum output power. The FS, DT, and NN pulses are all relatively robust against this. This is because many of the numerically shaped pulses use as much power as possible, so only the maximum power of the amplifier matters rather than the shape of the compression function at intermediate power. This is not entirely clear from the plots of the output pulse shapes in Fig. 4, but the input pulse shapes of $I$ and $Q$ plotted in Fig. S1 of the Supplementary Material sup are typically toggling between their maximum values of $-1$ and $1$, and the values plotted are only diminished from maximum amplitude due to the effects of the low-pass filter and resonator. Thus, changing the general shape of the amplifying function does not affect the pulse shapes much. The HS pulse shapes, on the other hand, do use the full range of the amplifier, and the dotted lines show how their performance is substantially diminished under the same $V_{\mathrm{sat}}$ reduction.
In the $B_{1}$ field attenuation plot in Fig. 5B, we plot the fractional reduction of the echo amplitude when the $B_{1}$ field is reduced to $80\%$. This probes how robust the echo amplitude is to $B_{1}$ field inhomogeneity. The best HS pulse is one of the more robust pulses against $B_{1}$ field inhomogeneity, since it is constructed as an adiabatic frequency sweep. Other pulse types include equivalently robust pulses, but generally there is no correlation between echo amplitude and robustness to inhomogeneity because the pulses were not optimized for robustness. In Fig. S3 of the Supplementary Material sup , we also plot echo amplitude versus resonator quality factor and versus $B_{1}$ field inhomogeneity.
@@
This ensemble of optimizations from different random seeds shows that there is no unique pulse shape that is optimal. Of course, there is always a trivial degeneracy of rotating both $I$ and $Q$ by ±$\pm 90^{\circ},\pm 180^{\circ}$ resulting only in a change of echo phase by the same angle (the degeneracy is four-fold rather than continuous due to the power constraint being a square rather than a circle in the I/Q plane), retaining the same refocusing time and echo amplitude. But beyond that, Fig. 5 shows that there are many pulse shapes with equivalent echo amplitudes that are not simply related by a phase transformation, as is clear from the different echo refocusing times. The fact that the landscape harbors many comparable maxima suggests that there remains a significant amount of flexibility in the parameterized pulses that could be used to satisfy additional constraints, such as favoring a particular refocusing time or phase (as in Kallies and Glaser (2018)), robustness to $B_{1}$ inhomogeneity, echo phase independence from $\tau_{1}$, or some other desirable property.
@@
Figure 6 strongly suggests that the improvement of the optimal FS, DT, and NN pulses compared to the HS pulse comes from i) improved performance near the band edges and ii) improved intra-band $\pi$ rotation performance in the presence of the power constraint. For a more comprehensive visualization of the spin dynamics, in Fig. S2 of the Supplementary Material sup we plot the total magnetization in $x$, $y$, and $z$ during each of the optimized pulse sequences.
@@
In other words, since all the spin packet phases are flipped by the $\pi$ rotation, echo formation requires $\tau_{\mathrm{eff},\pi/2}(\omega)-\tau_{\mathrm{eff},\pi}(\omega)$ to be a constant Kallies and Glaser (2018). Figure 7 shows that this is indeed the case for the FS, DT, and NN pulses, which have nonlinear phase dispersions for the two pulses but cooperate such that their nonlinear parts mutually cancel. Note that the difference $\tau_{\mathrm{eff},\pi/2}(\omega)-\tau_{\mathrm{eff},\pi}(\omega)=t_{\mathrm{refocus}}-\tau_{1}$. For example, in this case, the FS pulse has $\tau_{\mathrm{eff},\pi/2}(\omega)-\tau_{\mathrm{eff},\pi}(\omega)=$-10 ns and so refocuses at $90$ ns while the NN pulse has $\tau_{\mathrm{eff},\pi/2}(\omega)-\tau_{\mathrm{eff},\pi}(\omega)=39$ ns and refocuses at $139$ ns. The difference $\tau_{\mathrm{eff},\pi/2}(\omega)-\tau_{\mathrm{eff},\pi}(\omega)$ is shown in Fig. S4 of the Supplementary Material sup .
@@
These optimized pulse shapes have not yet been experimentally implemented. To record the echo generated by these pulses with sufficient fidelity, a receiver of sufficient bandwidth ($>$ 240 MHz) is required. Higher-bandwidth detection systems with up to 1 GHz of bandwidth have become available commercially recently (Bruker SpecJet 3 and VideoAmp 3), so the presented approach is timely and feasible. Also, non-commercial wideband receiver systems have been built Doll and Jeschke (2017).